# Multi-Excitation Projective Simulation with a Many-Body Physics Inspired Inductive Bias

### 15 Feb 2024 v1

@@ V.1 Invasion Game With Distraction
The Invasion Game is a standard toy environment [14] to visualize the basic concepts of PS. Mathematically, it is a special case of so-called contextual bandits problems [37, 38]. The original environment considered a sequence of doors that an attacker would try to enter through, and which a defender (the agent) would attempt to block for a set number of rounds. During each round, the attacker would indicate to the defender via some abstract symbols which door they intend to visit, and the defender would receive a reward based on whether they guessed the correct door or not. Thus, the task of the defender is to infer the meaning of said symbols by learning the attacker’s strategy. Here, we modify the Invasion Game such that it provides a simplistic environment showcasing the impact of different choices of few-body inductive biases.
@@ V.3 Computer Maintenance
The final environment we consider is that of diagnosing and fixing a broken computer, which we call the Computer Maintenance environment. Attempting to repair a computer is a fairly common task that many are familiar with. It can be quite a complex task with many possible causes in various systems giving rise to any particular problem, and yet, computer repair is not so complicated as to be completely intractable – as is evidenced by the multitude of successful computer repairs that occur daily. This containment of task complexity is mainly because the technicians who work on these problems can keep track of multiple variables simultaneously, which is a daunting and ultimately unfeasible situation for an agent who only considers single excitations at a time. It is for these reasons that we choose the Computer Maintenance environment to highlight the capabilities of an agent in a complex environment who considers multiple excitations throughout a chain-of-thought larger than length 1, and the usefulness of the inductive biases in cutting down the size of the ECM.
@@
We choose to visualize the Computer Maintenance environment as follows. A customer walks into a computer repair shop with their broken desktop computer and asks the technician if they can find the underlying cause(s) of the problem(s) and fix it(them). The technician agrees and then takes the broken computer back to their workshop to assess the situation. From this assessment, a number of symptoms indicative of possibly many underlying causes of the problem will become apparent to the technician, which could include a combination of software and hardware issues. It is then the technician’s task to identify the relevant components and assert a hypothesis about what the cause of the symptoms is given these components, then apply appropriate fixes that will hopefully solve the problem. Along with an explanation of the underlying cause of the problem, the technician also presents the customer with an invoice of how much time it took them to fix the problem and what the costs of the parts/software required were.
Translating the previous description to a reinforcement learning setting, the computer technician is interpreted as our mePS agent who can receive sets of symptoms generated from the environment as percepts and perform actions on the environment in the form of selecting sets of components and corresponding fixes to those components (as a pair of subsets). The environment contains a set of lists of the possible symptoms, components, causes, and fixes whose plain text descriptions are encoded as integers such that the agent is unaware of the association between the two a priori. The set of symptoms and their integer encodings used for this work are:
{‘PC overheating’: 1, ‘files disappearing’: 2, ‘visible markings on components’: 3, ‘unexpected shutdowns’: 4, ‘slow performance’: 5, ‘old hardware’: 6, ‘strange noises’: 7, ‘software glitches’: 8}.
The set of components and their integer encodings are:
{‘CPU’: 13, ‘SSD’: 14, ‘MoBo’: 15, ‘PSU’: 16, ‘OS’: 17}.
The set of causes and their integer encodings are:
{‘physical damage’: 18, ‘software damage’: 19, ‘malware’: 20, ‘faulty’: 21, ‘not connected’: 22}.
Finally, the set of fixes and their integer encodings are:
{‘replace components’: 9, ‘install missing software’: 10, ‘cooldown computer’: 11, ‘run antivirus’: 12}.
Elements from each of these sets are then combined to form what we call ‘scenarios,’ which are used to fix and specify a specific problem with a unique goal state that defines the length of a training episode: the agent applies their percept-to-action policy repeatedly until they reach the goal state, marking the end of the episode. At the beginning of each episode, a scenario is sampled uniformly at random and the corresponding subset of symptoms contained in it are then used for that episode. The chosen percept is fixed for the duration of the episode since this more closely corresponds with the real situation assuming no new problems arise during the repair process and that the agent has fully discovered all relevant symptoms beforehand such that they distinguish one problem from another. Allowing new symptoms to arise during each step would confuse the agent as to what the actual problem was since a given set of symptoms typically corresponds to only a small group of issues, thus preventing any explanation that agrees with the specified scenario.
@@
In this work, 11 different scenarios are being sampled that the agent must learn to navigate through and solve. To better demonstrate the savings accrued from reducing the size of the ECM, only scenarios involving at most 2 elements in any category are considered in the training set. For the sake of brevity, we only describe one of the scenarios: consider a clumsy owner who drops their computer on the ground, causing physical damage to the motherboard (MoBo) that results in a software error in the storage unit (SSD), leading to physical damage in the SSD, which then also causes physical damage to the MoBo in a feedback loop. Such a situation in which software interacts with hardware can occur when faulty programs overuse resources (resulting in heat or electrical damage to the computer) or create conflicting processes that lead to this. In the Computer Maintenance environment, this scenario would be coded as:
[[‘files disappearing’, ‘visible markings on components’], [‘MoBo’, ‘SSD’], [‘physical damage’, ‘software damage’], [‘replace components’]]
@@
It is immediately clear from this scenario that the feedback loop between the MoBo and the SSD demands that these components and the associated causes all be treated together if the agent hopes to solve the problem - something the mePS agent is better suited to handle. The only hope that the single-excitation agent has of solving the scenario is to exponentially inflate the size of their ECM so that all possible combinations of the environment variables are laid out in all of the clips. If each of the environment variables can take sufficiently many values, the previous strategy will undoubtedly fail since the random walk path in the agent’s ECM that represents the correct solution to a given scenario will have a vanishingly small probability of occurring.
@@
To implement longer chains-of-thought and highlight the benefit of moving from the single- to multi-excitation agent case, we incorporate a hidden layer into our mePS agent where the agent’s hypotheses about the underlying causes of the problem are represented by pairs of sets of components and causes. Therefore, we use two many-body $h$ values: the first for transitions between the percept layer and hidden layer, and the second for transitions between the hidden layer and action layer. This change to a 3-layer agent and the use of pairs of subsets as hidden and action layer elements introduces some modifications to how the inductive bias is applied to the mePS agent in this environment compared with the one described in the previous two environments. Firstly, because we have pairs of subsets for the hidden and action layer elements, greater flexibility than applying a many-body cutoff uniformly to the pair is important to ensure that the agent is not forced to consider a larger than-necessary subspace of the elements of the pairs with respect to the scenario key, whose elements can be subsets of unequal size in general. Not implementing this change would slow down learning unnecessarily, so we now assign a many-body cutoff value for each element in the hidden and action layer pairs. It is also necessary to have $IO$ contain all values up to and including a specified many-body cutoff value, for the same reason that the elements of the scenario key can be of different sizes. Lastly, because we have two many-body $h$ values now, we can specify different sets of many-body cutoffs for each of them.
@@
An inductive bias agent will be represented with the following notation: $\{n_{s},[n_{hc},n_{c}],[n_{ac},n_{f}]\}$, which we call an inductive bias configuration. Here, $n_{s}$ is the many-body cutoff on the number of symptoms in the percept layer represented in the inductive bias agent’s ECM; $n_{hc}$, the many-body cutoff on the number of components in the hidden layer; $n_{c}$, the many-body cutoff on the number of causes in the hidden layer; $n_{ac}$, the many-body cutoff on the number of components in the action layer; and $n_{f}$, the many-body cutoff on the number of fixes in the action layer. Since we are discarding leftover excitations, and because we have a layered, feed-forward architecture, the inductive bias used here corresponds with Inductive Bias 2DP. We can calculate the number of learning parameters $N_{l}$ for each agent from Equation (10) in Appendix A, which we will use throughout the rest of this section.
@@
Now that scenarios have been fully introduced, including the fact that scenarios are used by the environment as goal states, it is more natural to consider the technician’s explanation to the customer as the ‘action’ performed by the agent on the environment. Thus, the agent internally stores their ECM deliberation path and outputs it as an action on the environment. The environment then rewards the agent using two separate mechanisms, which we call the hypothesis and plausibility rewards, that are each applied to different $h$-values. We do this to avoid washing out the percept information that tends to happen when blindly applying the standard PS update rule to intermediate layers. The hypothesis reward is measured based on how close the chosen elements of the agent’s intermediate layer are to their corresponding partners in the given scenario; it is applied to the $h$-values between the percept and intermediate layers. If those elements match exactly, a reward of +5 is given, otherwise, a large penalty of -10 is applied. An additional component of the reward penalizes the agent by -1 if they pick more elements in both categories than those contained in the scenario key, and -0.5 if they only pick more elements in one of the categories; an attempt to discourage the agent from selecting the maximum number of elements permitted by their inductive bias (becomes more important for increasing many-body cutoff size).
@@
The plausibility reward, which is applied to the $h$-values between the intermediate layer and the action layer, has three different components to it. The first component has the same structure as the hypothesis reward but the values are quadrupled for the penalty on too many elements and it is for the agent’s chosen elements from the action layer instead of the intermediate layer. The second component checks whether the chosen components from the intermediate layer match those chosen from the action layer; a reward of +1 is given if they match perfectly, +0.25 if all of the action layer elements match but more intermediate layer elements are chosen then action layer elements, and -2 otherwise. This component of the reward can be interpreted as an internal consistency mechanism that encourages the agent to form coherent explanations between hypothesis and action. Note that it does not directly refer to the scenario key and is thus a general mechanism that can aid learning on scenarios not considered in the original training set. The third component of the reward is where the plausibility reward originally got its name from; it checks whether the agent’s chosen fixes make sense with respect to the underlying causes they identified in their explanation. This is judged based on certain causal relationships between what the causes and fixes each refer to, that are put in by hand. For example, implementing the fix ‘replace components’ would be justified if the cause of the problem was suspected to be ‘physical damage’ or ‘faulty,’ since both indicate problems with the hardware that are only fixable by physically removing them. However, if the suspected cause was simply ‘malware,’ then only selecting ‘replace components’ would not be justified because this is a fix one implements when there are hardware problems, not software. Now, for each possible fix there are specific causes that need to be selected for the agent to receive the associated reward of +0.3 divided by the number of fixes specified in the chosen scenario key $n_{\mathrm{fix}}$; this ensures that the agent can converge to the optimal reward since some scenarios require different numbers of fixes to solve than others, which would introduce oscillations about the optimal value if the agent got to that point in the training. A penalty of up to $-4/n_{\mathrm{fix}}$ is given if the proper causes are not selected. Aside from the use of $n_{\mathrm{fix}}$ (which could be replaced with the number of fixes chosen by the agent, in retrospect), the third component of the plausibility reward is also independent of the scenario key, adding another general mechanism for augmenting learning beyond the training set.
@@
Finally, if all elements in the agent’s explanation exactly match those in the scenario key, the agent receives a big bonus to both rewards of +15 to amplify the probability of selecting this deliberation path again in future episodes. Once this bonus is triggered it signals an end to the episode and a fresh set of symptoms corresponding to another scenario are then selected. Before receiving this signal and before the agent factors in the reward into their update rule definition 6, the reward is first sent to an external reward shaping function defined by Equation (11) whose purpose is to discourage the agent from taking too many steps within an episode, with the aim of curtailing arbitrarily long episodes; details about this function can be found in Appendix A.
@@
All agents for each of their layers use the softmax function with $\beta=1/2$ to convert $h$-values to probabilities. Additionally, all $h$-values are initialized to 1. The results of the training are shown in Figure 6. We can see that the inductive bias agent achieves near-optimal values for both the hypothesis and plausibility rewards using far fewer steps but requiring more episodes than the unrestricted agent does, as seen in the total step number per episode curve, which better illustrates the savings from the inductive bias as the two average reward plots do not show how many total steps have elapsed during training. To quantify the difference in step number to the optimal reward, we calculated the total average steps taken over 200 episodes for both agents: the inductive bias agent takes roughly 5474 steps while the unrestricted agent takes roughly 8168 steps. Together with the complexity-theoretic results from Section IV.3, we only expect this difference in step number to grow. Although the unrestricted agent does converge to the optimal rewards in fewer episodes, it clearly takes more steps to do so because it has roughly 16 times the number of trainable parameters as the inductive bias agent has. The fact that we get comparable performance from the inductive bias agent using far fewer parameters justifies the relatively small fluctuations around the optimal value for each of the reward curves, which are expected to decrease as more agents are included in the average. The total step number per episode curve also shows roughly exponential decay from the agent trying out scenario configurations randomly to always picking the right configuration immediately.

### 29 Feb 2024 v2

@@ V.1 Invasion Game With Distraction
Figure 4: The defender $D$ must guess which door the attacker $A$ will go to based on a set of symbols shown to them, and block $A$. $D$ is rewarded for success and punished for failure.
The Invasion Game is a standard toy environment [14] to visualize the basic concepts of PS. Mathematically, it is a special case of so-called contextual bandits problems [37, 38]. The original environment considered a sequence of doors that an attacker would try to enter through, and which a defender (the agent) would attempt to block for a set number of rounds. During each round, the attacker would indicate to the defender via some abstract symbols which door they intend to visit, and the defender would receive a reward based on whether they guessed the correct door or not (as visualized in Figure 4). Thus, the task of the defender is to infer the meaning of said symbols by learning the attacker’s strategy. Here, we modify the Invasion Game such that it provides a simplistic environment showcasing the impact of different choices of few-body inductive biases.
@@ V.3 Computer Maintenance
The final environment we consider is that of diagnosing and fixing a broken computer, which we call the Computer Maintenance environment. Computer repair is an everyday task that can be quite complex, with many possible causes in various systems giving rise to any particular problem; and yet, the task is accomplished daily, making it not so complicated as to be completely intractable. Computer technicians can manage this task complexity because they can keep track of multiple variables simultaneously, which is a daunting and ultimately unfeasible situation for an agent restricted to single excitations only. It is for these reasons that we choose the Computer Maintenance environment to highlight the capabilities of an agent in a complex environment who considers multiple excitations throughout a non-trivial chain-of-thought, as well as the usefulness of the inductive biases in reducing ECM size.
@@
We choose to visualize the Computer Maintenance environment as follows. A customer enters a computer repair shop with a broken computer and asks the technician to find the root causes of the issues and fix them. While the technician is diagnosing the problem, several symptoms indicative of possibly many underlying causes of the problem become apparent, which could include a combination of software and hardware issues. The technician must now identify the relevant components and assert a hypothesis about what the causes of the symptoms are given these components, then apply appropriate fixes that will hopefully solve the problem. Along with an explanation of the underlying causes of the problem, the technician also presents an invoice to the customer that states the amount of time and resources used to repair the computer.
@@
Figure 7: A schematic illustration of one environment step and the mePS agent’s ECM during this step (in the Computer Maintenance environment). The symptoms are represented by the blue dots; the components by grey dots; the causes by orange dots; and the fixes by red dots (the numbers correspond to the place of that object in Table 1). The boxes denote pairs of excitation configurations that the agent can transition to/from. The agent first observes a pair of symptoms that get coupled into a percept source hyperedge $S_{1}$ which triggers deliberation through pairs of components/causes ($C_{i},C^{{}^{\prime}}_{j}$) and components/fixes ($C_{i},A_{j}$), ending with a chain-of-thought that the agent then converts into an explanation to the customer who determines their reward. As an example, consider a clumsy owner who drops their computer on the ground, causing physical damage to the motherboard (MoBo) that results in a software error in the storage unit (SSD), leading to physical damage in the SSD, which then also causes physical damage to the MoBo in a feedback loop. Such a situation in which software interacts with hardware can occur when faulty programs overuse resources (resulting in heat or electrical damage to the computer) or create conflicting processes that lead to this. In the Computer Maintenance environment, this scenario would be coded as: [[‘files disappearing’, ‘visible markings on components’], [‘MoBo’, ‘SSD’], [‘physical damage’, ‘software damage’], [‘replace components’]].
Translating the previous description to a reinforcement learning setting, the computer technician is interpreted as our mePS agent who can receive sets of symptoms generated from the environment as percepts and perform actions on the environment in the form of selecting sets of components and corresponding fixes to those components (as a pair of subsets). The environment contains a set of lists of the possible symptoms, components, causes, and fixes whose text descriptions are encoded as integers such that the agent is unaware of the association between the two a priori; Table 1 displays them.
Symptoms
Components
Causes
Fixes
PC overheating: 1
CPU: 13
physical damage: 18
replace components: 9
files disappearing: 2
SSD: 14
software damage: 19
install missing software: 10
visible markings on components: 3
MoBo: 15
malware: 20
cooldown computer: 11
unexpected shutdowns: 4
PSU: 16
faulty: 21
run antivirus: 12
slow performance: 5
OS: 17
not connected: 22
old hardware: 6
strange noises: 7
software glitches: 8
Table 1: The symptoms, components, causes, and fixes in the Computer Maintenance environment along with their integer encodings.
Elements from each of these sets are then combined to form what we call ‘scenarios,’ which are used to fix and specify a specific problem with a unique goal state that defines the length of a training episode: the agent repeatedly applies their policy until reaching the goal state, marking the end of the episode. At the beginning of each episode, a scenario is sampled uniformly at random and the corresponding subset of symptoms contained in it are then used for that episode. The chosen percept is fixed for the duration of the episode to more closely emulate the real situation assuming no new problems arise during the repair process and that the symptom set distinguishes one problem from another. Allowing new symptoms to arise during each step would confuse the agent on what the actual problem was since a given symptom set typically only corresponds to a small group of issues, thus preventing any explanation that agrees with the specified scenario.
@@
The agent must learn to navigate and solve 11 different scenarios in this work, which we constrained to have at most 2 elements per category to better demonstrate the savings accrued from reducing the size of the ECM. An example scenario is detailed in Figure 7. It is immediately clear from this example scenario that the feedback loop between the MoBo and the SSD demands that these components and the associated causes all be treated together if the agent hopes to solve the problem – something the mePS agent is better suited to handle. The only hope that the single-excitation agent has of solving the scenario is to exponentially inflate the size of their ECM so that all possible combinations of the environment variables are distributed in all clips. If each environment variable can take sufficiently many values, the previous strategy will undoubtedly fail since the random walk path in the agent’s ECM that represents the correct solution to a given scenario will have a vanishingly small probability of occurring.
@@
To implement longer chains-of-thought and highlight the benefit of moving from the single- to multi-excitation agent case, we incorporate a hidden layer into our mePS agent where the agent’s hypotheses about the underlying causes of the problem are represented by pairs of sets of components and causes. Therefore, we use two many-body $h$ values: the first for transitions between the percept layer and hidden layer, and the second for transitions between the hidden layer and action layer. This change to a 3-layer agent and the use of pairs of subsets as hidden and action layer elements introduces some modifications to how the inductive bias is applied compared to the previous two Invasion Games. Firstly, because we have pairs of subsets for some layers, a many-body cutoff applied to the pair will sometimes force the agent to consider a subspace for one of the subsets that is larger than the scenario demands, since the scenario elements can be subsets of unequal size in general; this slows down learning unnecessarily, so many-body cutoff values are now assigned for each element in the pairs. It is also necessary to have $IO$ contain all values up to and including a specified many-body cutoff value, for the same reason stated in the previous sentence. Lastly, because we have two many-body $h$ values now, we can specify different sets of many-body cutoffs for each of them.
@@
An inductive bias agent will be represented with the following notation: $\{n_{s},[n_{hc},n_{c}],[n_{ac},n_{f}]\}$, which we call an inductive bias configuration. Here, $n_{s}$ is the many-body cutoff on the number of symptoms in the inductive bias agent’s percept layer; $n_{hc}$, the cutoff on the number of components in the hidden layer; $n_{c}$, the cutoff on the number of causes in the hidden layer; $n_{ac}$, the cutoff on the number of components in the action layer; and $n_{f}$, the cutoff on the number of fixes in the action layer. Since we are discarding leftover excitations, and because we have a layered, feed-forward architecture, the inductive bias used here corresponds with Inductive Bias 2DP. We can calculate the number of learning parameters $N_{l}$ for each agent from Equation (10) in Appendix A, which we will use throughout the rest of this section.
@@
Figure 7 illustrates the process leading up to the agent receiving a reward from the environment. The reward is decomposed into two separate mechanisms that are each applied to different $h$-values, which we call the hypothesis and plausibility rewards. We do this to avoid washing out the percept information that tends to happen when blindly applying the standard PS update rule to intermediate layers. The hypothesis reward is measured based on how close each intermediate layer element is to its corresponding partner in the given scenario; it is applied to the $h$-values between the percept and intermediate layers. If those elements match exactly, a reward of +5 is given, otherwise, a large penalty of -10 is applied. An additional component of the reward penalizes the agent by up to -1 if they pick more elements than there are in the given scenario; an attempt to discourage the agent from selecting the maximum number of elements permitted by their inductive bias (becomes more important for increasing many-body cutoff size).
@@
The plausibility reward, which is applied to the $h$-values between the intermediate layer and the action layer, has three different components to it. The first component has the same structure as the hypothesis reward but the values are quadrupled for the penalty on too many elements and it is for the agent’s chosen elements from the action layer instead of the intermediate layer. The second component checks whether the chosen components from the intermediate layer match those chosen from the action layer; a reward of +1 is given if they match perfectly, +0.25 if all of the action layer elements match but more intermediate layer elements are chosen then action layer elements, and -2 otherwise. This component of the reward can be interpreted as an internal consistency mechanism that encourages the agent to form coherent explanations between hypothesis and action. Note that it does not directly refer to the scenario key and is thus a general mechanism that can aid learning on scenarios not considered in the original training set. The third component of the reward checks whether the agent’s chosen fixes make sense with respect to the underlying causes they identified in their explanation. This is judged based on certain causal relationships between what the causes and fixes each refer to, that are put in by hand. For example, implementing the fix ‘replace components’ would be justified if the cause of the problem was suspected to be ‘physical damage’ or ‘faulty,’ since both indicate problems with the hardware that are only fixable by physically removing them. However, if the suspected cause was simply ‘malware,’ then only selecting ‘replace components’ would not be justified because this is a fix one implements when there are hardware problems, not software. Now, for each possible fix there are specific causes that need to be selected for the agent to receive the associated reward of +0.3 divided by the number of fixes specified in the chosen scenario key $n_{\mathrm{fix}}$; this ensures that the agent can converge to the optimal reward since some scenarios require different numbers of fixes to solve than others. A penalty of up to $-4/n_{\mathrm{fix}}$ is given if the proper causes are not selected. Aside from the use of $n_{\mathrm{fix}}$, the third component of the plausibility reward is also independent of the scenario key, adding another general mechanism for augmenting learning beyond the training set.
@@
Finally, if all elements in the agent’s explanation exactly match those in the scenario key, the agent receives a big bonus to both rewards of +15 to amplify the probability of selecting this deliberation path again in future episodes. Once this bonus is triggered it signals an end to the episode and a fresh set of symptoms corresponding to another scenario are then selected. Before receiving this signal and before the agent factors in the reward into their update rule definition 6, the reward is first sent to an external reward shaping function defined by Equation (11) whose purpose is to discourage the agent from taking too many steps within an episode, intending to curtail arbitrarily long episodes; details about this function can be found in Appendix A.
@@
All agents for each of their layers use the softmax function with $\beta=1/2$ to convert $h$-values to probabilities and all $h$-values are initialized to 1. The results of the training are shown in Figure 8. We can see that the inductive bias agent achieves near-optimal values for both the hypothesis and plausibility rewards using far fewer steps but requiring more episodes than the unrestricted agent does, as seen in the total step number per episode curve, which better illustrates the savings from the inductive bias as the two average reward plots do not show how many total steps have elapsed during training. The total step number per episode curve also shows roughly exponential decay, which is indicative of the agent first trying out scenario configurations randomly and then always picking the right configuration immediately. To quantify the difference in step number to the optimal reward, we calculated the total average steps taken over 200 episodes for both agents: the inductive bias agent takes roughly 5474 steps while the unrestricted agent takes roughly 8168 steps. Together with the complexity-theoretic results from Section IV.3, we only expect this difference in step number to grow. Although the unrestricted agent does converge to the optimal rewards in fewer episodes, it clearly takes more steps to do so because it has roughly 16 times the number of trainable parameters as the inductive bias agent has. The fact that we get comparable performance from the inductive bias agent using far fewer parameters justifies the relatively small fluctuations around the optimal value for each of the reward curves, which are expected to decrease as more agents are included in the average.